# Can Large Language Models Understand As Well As Apply Patent Regulations to Pass a Hands-On Patent Attorney Test?

### v1

@@ Abstract
OpenAI o1 led with $0.82$ accuracy and $0.81$ F1 score, whereas (Amazon Web Services) AWS Llama 3.1 8B lagged at $0.50$ accuracy, and a Python-deployed Llama 3.1 8B scored $0.55$. The latter two are within the range of mere guessing for the two-answer forced-choice design. None of the evaluated models could have passed the examination fully, as accuracy never exceeded the average threshold of $0.90$ required for professional-level standards—also not models that are regularly promoted for their assumed beyond-PhD- and bar-admitted-lawyer-level performance. GPT-4o excelled at integrating text and graphics, while Claude 3 Opus often lost formatting coherence.
Human patent experts evaluated the textual justifications and uncovered various critical shortcomings of each model. They valued clarity and legal rationale over the raw correctness of the answers, which revealed misalignment between automatic metrics and expert judgment. Model outputs were sensitive to modest temperature changes and prompt wording, which underscores the remaining necessity of expert oversight. Future work should target logical consistency, robust multimodality, and adaptive prompting to approach human-level patent proficiency. In summary, despite the outstanding performance of recent large models, the general public might overestimate their performance. The field has a long way to go to develop a virtual patent attorney. This paper wants to point out several specific limitations that need solutions.
@@ 4 Results
Figure 3: Accuracy of each LLM on the EQE pre-exam; blue bars show per-model accuracy, the red dashed line marks random guessing (0.5), and the pink band indicates the ±$\pm 1/\sqrt{n}$ fluctuation range discussed above.
@@ 4.3 Context-Length Management
[…]
@@ 4.6 Year- and Question-Specific Performance
Across the twelve EQE pre-exams (2012–2024, in Appendix A), the panel-wide mean accuracy climbed from $0.59$ in 2012 to a peak of $0.77$ in 2013, after which it fluctuated between $0.58$ and $0.69$ (median $0.63$). OpenAI o1 set the yearly benchmark accuracy in ten of the twelve papers and averaged $0.82\pm 0.07$, while Sonnet 3.5 briefly excelled in 2018 ($0.93$ mean accuracy in 2018) and again in 2022 ($0.83$ mean accuracy in 2022). When you look at the accuracy for each individual statement (not just overall or by year), you can see that the model’s success depends more on the specific content or challenge of each question than on the order in time (chronology / year) in which the questions appeared. For example, in the 2024 paper, the four nested statements of Question 1 were solved by none of the thirteen models correctly, whereas only three models could answer all the statements from Question 5 correctly. OpenAI o1 was perfect on Questions $6$, $7$, $9$, and $10$. However, for the first cluster (from Questions $1$ to $5$), OpenAI o1 was right only for question $4$. Hence, there were huge differences in the accuracies of both the clusters. For the first cluster, it was $0.75$ and for the second was $0.95$, hence the total mean accuracy in 2024 for OpenAI was $0.85$. This kind of uneven trend, struggling early, then excelling later, was also observed in Sonnet 3, but not in the other models.
@@ 4.7.2 General Observations Noted by Human Experts
The Python implementation of Llama 3.1 8B as one of the smallest models in the comparison attempted to include citations—sometimes in large numbers—in various formats (e.g., just reference numbers in square brackets, sometimes with, sometimes without a corresponding bibliography, specific rule numbers, website addresses, etc.). However, most of these references were either fabricated or incorrect, particularly those referring to the European Patent Convention (EPC) and the Patent Cooperation Treaty (PCT). It failed to distinguish between novelty and obviousness and was unable to interpret dates correctly. It sometimes calculated the correct deadline in the explanation, but still disagreed with the correct outcome, which evaluators described as “absurd”. Additionally, the model’s explanations were often overly verbose, repetitive, and incorrect, with issues in coherence and legal soundness. The model from time to time referred to specific requirements in EPC articles, which indicates that the respective or related texts were in the training set but potentially overwhelmed the model structure and flexibility.
@@ 4.8 Comparative Analysis of GPT-4o and Opus
The professional human evaluation provided numerical scores (from 1 to 10 linearly scaled to 0 to 1 in postprocessing) on individual questions for conceptual clarity, coherence, and justification quality beyond mere correctness. Upon closer analysis, it was observed that models such as Sonnet 3.5 and Python Llama 3.1 8B negatively correlated with correctness (respectively $-0.07$ and $-0.16$). Thus, evaluators awarded scores more for justification quality and logical consistency rather than solely on the True or False answers so that random guessing could not score points easily, and correct reasoning, fact extraction from the questions, and subsumption were necessary. This observation underscores the nuanced nature of human evaluation and highlights the critical role that qualitative factors play in expert decision-making processes. Thus, human oversight remains essential, as computational accuracy alone does not fully encapsulate expert-level understanding or reasoning.

### v2

@@ Abstract
OpenAI o1 led with $0.82$ accuracy and $0.81$ F1 score, whereas (Amazon Web Services) AWS Llama 3.1 8B lagged behind at $0.50$ accuracy, and a Python-deployed Llama 3.1 8B scored $0.55$. The latter two are within the range of mere guessing for the two-answer forced-choice design. None of the evaluated models could have passed the examination fully, as accuracy never exceeded the average threshold of $0.90$ required for professional-level standards—also not models that are regularly promoted for their assumed beyond-PhD- and bar-admitted-lawyer-level performance. GPT-4o excelled at integrating text and graphics, while Claude 3 Opus often lost formatting coherence.
Human patent experts evaluated the textual justifications and uncovered various characteristic traits as well as critical shortcomings of each model. They valued clarity and legal rationale over the raw correctness of the answers, which revealed misalignment between automatic metrics and expert judgment. Model outputs were sensitive to modest temperature changes and prompt wording, which underscores the remaining necessity of expert oversight. Future work should target logical consistency, robust multimodality, and adaptive prompting to approach human-level patent proficiency. In summary, despite the outstanding performance of recent large models, the general public might overestimate their performance. The field has a long way to go to develop a virtual patent attorney. This paper wants to point out several specific limitations that need solutions.
@@ 4 Results
Figure 3: Accuracy of each LLM on the EQE pre-exam; blue bars show per-model accuracy, the red dashed line marks random guessing (0.5), and the pink band indicates the $\pm 1/\sqrt{n}$ fluctuation range discussed above.
@@ 4.3 Context-Length Management
For every other model—Meta Llama 3.1 405B, Meta Llama 3.1 70B, Meta Llama 3.1 8B, Mistral Large, Anthropic Claude 3 Sonnet, Anthropic Claude 3.5 Sonnet, OpenAI GPT-4o, Deepseek-v3, and Deepseek-v3 R1—we adopted the more demanding setting: the first ten legal questions were concatenated into a single prompt ( $\approx\,1\,700$ tokens). This longer prompt length was obtained by summing each question’s token count after running them all through the same tokenizer, mirroring how a practitioner might supply an entire section of the EQE in one API call.
The multimodal trial increased the level of difficulty: the complete PDF of Parts 1–3 (20 questions plus claim drawings,  $\approx\,14\,000$ tokens after optical character recognition with figure captions kept) was passed in one shot to OpenAI GPT-4o and Anthropic Claude 3 Opus. Under these conditions, GPT-4o maintained $0.70$ accuracy, whereas Opus dropped to $0.67$. Thus, the various transformer models and their self-attention designs appear to differ more in cache policy and long-context fine-tuning than in raw parameter count [16].
@@ 4.6 Year- and Question-Specific Performance
Across the twelve EQE pre-exams (2012–2024, in Appendix A), the panel-wide mean accuracy climbed from $0.59$ in 2012 to a peak of $0.77$ in 2013, after which it fluctuated between $0.58$ and $0.69$ (median $0.63$). OpenAI o1 set the yearly benchmark accuracy in ten of the twelve papers and averaged $0.82\pm 0.07$, while Sonnet 3.5 briefly excelled in 2018 ($0.93$ mean accuracy in 2018) and again in 2022 ($0.83$ mean accuracy in 2022). When you look at the accuracy for each individual statement (not just overall or by year), you can see that the model’s success depends more on the specific content or challenge of each question than on the order in time (chronology / year) in which the questions appeared. For example, in the 2024 paper, the four nested statements of Question 1 were solved by none of the thirteen models correctly, whereas only three models could answer all the statements from Question 5 correctly. OpenAI o1 was perfect on Questions $6$, 77, 99, and $10$. However, for the first cluster (from Questions $1$ to $5$), OpenAI o1 was right only for question $4$. Hence, there were huge differences in the accuracies of both the clusters. For the first cluster, it was $0.75$ and for the second was $0.95$, hence the total mean accuracy in 2024 for OpenAI was $0.85$. This kind of uneven trend, struggling early, then excelling later, was also observed in Sonnet 3, but not in the other models.
@@ 4.7.2 General Observations Noted by Human Experts
The Python implementation of Llama 3.1 8B as one of the smallest models in the comparison attempted to include citations—sometimes in large numbers—in various formats (e.g., just reference numbers in square brackets, sometimes with, sometimes without a corresponding bibliography, specific rule numbers, website addresses, etc.). However, most of these references were either fabricated or incorrect, particularly those referring to the European Patent Convention (EPC) and the Patent Cooperation Treaty (PCT). It failed to distinguish between novelty and obviousness and was unable to interpret dates correctly. It sometimes calculated the correct deadline in the explanation, but still disagreed with the correct outcome, which evaluators described as “absurd”. Additionally, the model’s explanations were often overly verbose, repetitive, and incorrect, with issues in coherence and legal soundness. The model from time to time referred to specific requirements in EPC articles, which indicates that the respective or related texts were in the training set but potentially overwhelmed the model structure and flexibility
More critical appears to be that the model sometimes hallucinated wrong facts and added them to the case (including dates that were not given).
@@
DeepSeek v3
DeepSeek v3 practically never cited any regulations. It struggled from time to time with sentences that contained many conditions in if–then logics and could either not subsume correctly or missed the logical connection of the requirements (e.g., mandatory to fulfill all vs. only one). In various cases, the questions—which were often phrased as statements—appeared to be taken by the model as fact so that they replaced regulation facts. Mistral appeared to demonstrate a similar behavior.
More importantly, DeepSeek v3 missed more intricate novelty–inventiveness rules in combination with priorities and / or not yet published prior patent applications. The model, for instance, missed the point that patent applications have to be novel over earlier filed but not yet published earlier applications but not inventive. DeepSeek v3 however could apply the novelty concept well and differentiate it from novelty if confronted very specifically with combinations of feature sets that were novel but not necessarily inventive. It demonstrated similar weaknesses in regulations about divisional applications, which it shared with other models (e.g., GPT4o, Gemma 2, or Llama 8B), although some models (e.g., Opus, Mistral Large, Llama 70B, OpenAI o1, and also DeepSeek R1) managed them. However, it was not clear if wrong answers and explanations were caused by knowledge gaps, i.e., missing information in the training data, or an issue with language logic. The latter would align with the observation of failure to deal with more intricate logic statements. However, DeepSeek v3 certainly had exposure to PCT and / or EPC regulations or secondary literature about them in its training repertoire as it could answer some pure knowledge questions that were too specific to hallucinate them correctly or take them from other jurisdictions.
The model also struggled with deadline calculations when at least one time was only given roughly (e.g., month or season names) instead of an exact date. Except for Llama 8B, most other models appeared to deal better with such cases. DeepSeek v3 also miscalculated deadlines where the question provided exact dates.
Overall, the explanations were typically very brief and compact, which impeded a further analysis. In a few cases, such single-sentence answers were very close to the short answers provided by the EPO, while widely wrong answers in many other questions would not support the hypothesis that the entire set of brief answers was in the training set.
DeepSeek R1
The DeepSeek reasoning model performed very differently according to the experts’ observations. The model still does not often cite references but when it does, they were mostly right. The answers are not as recitent as v3 but sometimes delivered page-long, highly colloquial (“This is tricky”, “Wait, no”) texts that between dialogues and streams of consciousness. With the information not given to the evaluators that these answers belonged to a reasoning model, clarifies that many answers appear to present the often hard to read raw output stream in which statements are iteratively processed and combined. The model from time to time speculated and hypothetically expanded cases but in contrast to hallucinations of other models clearly stated those aspects subjunctively.
It could answer and correctly explain several cases, where DeepSeek v3 failed. Specifically, as the patent experts commented on it in both cases, the regulations around European divisional applications, their priorities, and the limitation to the original disclosure.
The calculation skills of deadlines and fees were mostly right, but missed holidays. DeepSeek R1 often provided justification or breakdowns for calculations that allowed a deeper look and in various cases accounted for special cases, such as delivery aspects (e.g., Rule 126 EPC), though not always unambiguously correct. It missed holidays though and has shown cases where it speculated in subjunctive form that a date might fall onto a weekend, which could move the deadline to the next business day instead of knowing it for the specific date.
@@ 4.8 Comparative Analysis of GPT-4o and Opus
The professional human evaluation provided numerical scores (from 1 to 10 linearly scaled to 0 to 1 in postprocessing) on individual questions for conceptual clarity, coherence, and justification quality beyond mere correctness. Upon closer analysis, it was observed that models such as Sonnet 3.5 and Python Llama 3.1 8B negatively correlated with correctness (respectively 0.07-0.07 and 0.16-0.16). Thus, evaluators awarded scores more for justification quality and logical consistency rather than solely on the True or False answers so that random guessing could not score points easily, and correct reasoning, fact extraction from the questions, and subsumption were necessary. This observation underscores the nuanced nature of human evaluation and highlights the critical role that qualitative factors play in expert decision-making processes. Thus, human oversight remains essential, as computational accuracy alone does not fully encapsulate expert-level understanding or reasoning.